# Click-Calib: A Robust Extrinsic Calibration Method for Surround-View Systems

### v1

@@ Abstract
Surround-View System (SVS) is an essential component in Advanced Driver Assistance System (ADAS) and requires precise calibrations. However, conventional offline extrinsic calibration methods are cumbersome and time-consuming as they rely heavily on physical patterns. Additionally, these methods primarily focus on short-range areas surrounding the vehicle, resulting in lower calibration quality in more distant zones. To address these limitations, we propose Click-Calib, a pattern-free approach for offline SVS extrinsic calibration. Without requiring any special setup, the user only needs to click a few keypoints on the ground in natural scenes. Unlike other offline calibration approaches, Click-Calib optimizes camera poses over a wide range by minimizing reprojection distance errors of keypoints, thereby achieving accurate calibrations at both short and long distances. Furthermore, Click-Calib supports both single-frame and multiple-frame modes, with the latter offering even better results. Evaluations on our in-house dataset and the public WoodScape dataset demonstrate its superior accuracy and robustness compared to baseline methods. Code is avalaible at https://github.com/lwangvaleo/click_calib.
@@ 3.2 Fisheye camera models
where $\theta$ denotes the incident angle and $r$ represents image radius in pixels. The coefficients $a_{1}$ to $a_{4}$ are distortion parameters distortion parameters obtained from intrinsic calibration. Fig. 4 illustrates the fisheye projection and compares it with the pinhole camera model.
Figure 4: Fisheye camera model. $\mathbf{P}^{C}$ is a 3D point in camera coordinate system. It intersects the unit sphere at $\mathbf{S}^{C}$ and is projected to $\mathbf{p}$ in the fisheye image, whereas it would be projected to $\mathbf{p^{\prime}}$ by a pinhole camera.
@@
$\alpha=\arctan\left(\frac{v-\Delta{v}}{u-\Delta{u}}\right)$ (4)
@@
$r=\sqrt{(u-\Delta{u})^{2}+(v-\Delta{v})^{2}}$ (5)
$\Delta{u}$ and $\Delta{v}$ are the pixel coordinates of the image center $O$.
@@ 3.3 Camera-vehicle projection
where $\mathbf{P}_{g}^{V}$ is the ground point in $V$. Here, $Z_{g}^{V}=0$ because the origin of the vehicle coordinate is located on the ground. With this constraint, the sclaing factor $\lambda$ can be uniquely determined, and $\mathbf{P}_{g}^{V}$ can then be calculated from Eq. 12.
@@ 4.1 Experimental setup
The proposed Click-Calib are tested on two datasets: our in-house dataset (collected by two cars, referred to as Car 1 and Car 2 in the following sections) and the public WoodScape dataset[31] (collected by one car). All three cars are equipped with four fisheye SVS cameras, providing 360° coverage around the ego vehicle with overlapping zones, as illustrated in Fig. 2. The image resolution in our dataset is is 1280 × 800, while in the WoodScape dataset it is 1280 × 966. To avoid image desynchronization issues at high speed, we only consider frames where the car’s speed is less than 30 km/h. The collected images cover three critical scenarios—indoor parking, outdoor parking, and city driving—to demonstrate the robustness of the proposed approach.

### v3

@@ Abstract
Surround-View System (SVS) is an essential component in Advanced Driver Assistance System (ADAS) and requires precise calibrations. However, conventional offline extrinsic calibration methods are cumbersome and time-consuming as they rely heavily on physical patterns. Additionally, these methods primarily focus on short-range areas surrounding the vehicle, resulting in lower calibration quality in more distant zones. To address these limitations, we propose Click-Calib, a pattern-free approach for offline SVS extrinsic calibration. Without requiring any special setup, the user only needs to click a few keypoints on the ground in natural scenes. Unlike other offline calibration approaches, Click-Calib optimizes camera poses over a wide range by minimizing reprojection distance errors of keypoints, thereby achieving accurate calibrations at both short and long distances. Furthermore, Click-Calib supports both single-frame and multiple-frame modes, with the latter offering even better results. Evaluations on our in-house dataset and the public WoodScape dataset demonstrate its superior accuracy and robustness compared to baseline methods. Code is available at https://github.com/lwangvaleo/click_calib.
@@ 3.2 Fisheye camera models
where $\theta$ denotes the incident angle and $r$ represents image radius in pixels. The coefficients $a_{1}$ to $a_{4}$ are distortion parameters obtained from intrinsic calibration. Fig. 4 illustrates the fisheye projection and compares it with the pinhole camera model.
Figure 4: Fisheye camera model. $\mathbf{P}^{C}$ is a 3D point in camera coordinate system. It intersects the unit sphere at $\mathbf{S}^{C}$ and is projected to $\mathbf{p}$ in the fisheye image (blue ray), whereas it would be projected to $\mathbf{p^{\prime}}$ by a pinhole camera (black ray).
@@
$\alpha=\arctan\left(\frac{v-v_{0}}{u-u_{0}}\right)$ (4)
@@
$r=\sqrt{(u-u_{0})^{2}+(v-v_{0})^{2}}$ (5)
where $(u_{0},v_{0})$ denotes the principal point in pixel coordinates.
@@ 3.3 Camera-vehicle projection
where $\mathbf{P}_{g}^{V}$ is the ground point in $V$. Here, $Z_{g}^{V}=0$ because the origin of the vehicle coordinate is located on the ground. With this constraint, the scaling factor $\lambda$ can be uniquely determined, and $\mathbf{P}_{g}^{V}$ can then be calculated from Eq. 12.
@@ 4.1 Experimental setup
The proposed Click-Calib is tested on two datasets: our in-house dataset (collected by two cars, referred to as Car 1 and Car 2 in the following sections) and the public WoodScape dataset[31] (collected by one car). All three cars are equipped with four fisheye SVS cameras, providing 360° coverage around the ego vehicle with overlapping zones, as illustrated in Fig. 2. The image resolution in our dataset is is 1280 × 800, while in the WoodScape dataset it is 1280 × 966. To avoid image desynchronization issues at high speed, we only consider frames where the car’s speed is less than 30 km/h. The collected images cover three critical scenarios—indoor parking, outdoor parking, and city driving—to demonstrate the robustness of the proposed approach.